Frank can do in 8 days a task which he and James together can do in three days. How many days would be required for James to do the work alone?
Answer by richwmiller(17219)

rate for A alone 8
rate for A & B together 3
1/a+1/b=1/x
1/8+1/b=1/3
1/b=1/3-1/8
1/b=8/24-3/24
1/b=5/24
1/b=1/4.8
b=4.8
rate for B alone 4.8
check
1/8+1/4.8=1/3
4.8/38.4+8/38.4=1/3
12.8/38.4=1/3
ok
codewa